OpenStudy (anonymous):

Find the diameter of a cone that has a volume of 56.52 cubic inches and a height of 6 inches. Use 3.14 for pi. a. 3 inches b. 6 inches c. 9 inches d. 12 inches

OpenStudy (yttrium):

Do you know volume of a cone's formula?

Directrix (directrix):

Cone Volume Formula (Attached)

Directrix (directrix):

@lowcard2

OpenStudy (wolf1728):

Volume = 56.52 height =6 Volume = 1/3 PI * r² * h 56.52 = 1/3 * PI * r² * 6 Now solve for 'r'

OpenStudy (anonymous):

a. 3inches 56.52/3.14=18 18/6=3 3*(3/1)= 9 9square rooted = 3